# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v5

@@ Abstract.
We formulate a precise connection between the new Drinfeld presentation of a quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$ and the new Drinfeld presentation of affine coideal subalgebras of split type recently discovered by Lu and Wang. In particular, we establish a “factorization formula”, expressing the commuting “Drinfeld–Cartan”-type operators $\Theta_{i,k}$ in a coideal subalgebra in terms of the corresponding Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$. We study the spectra of these operators on finite dimensional representations, and describe them in terms of rational functions with an extra symmetry. These results can be seen as the starting point of a $q$-character theory for affine Kac–Moody quantum symmetric pairs. Additionally, we prove a compatibility result linking Lusztig’s and the Lu-Wang-Zhang braid group actions.
@@ 1.1. A trinity of presentations
It is well known that quantum affine algebras $U_{q}\widehat{{\mathfrak{g}}}$ admit three distinct presentations: the Kac–Moody (also: Drinfeld–Jimbo, Serre) presentation [Jim85, Dri87b], the new Drinfeld (also: current or loop) presentation [Dri87a, Bec94, Dam12, Dam15] and the FRT (also: RTT) presentation [RSTS90]. The first of these can be seen as a quantization of the usual presentation of an affine Kac–Moody Lie algebra, while the second quantizes its construction as a central extension of a loop Lie algebra. Precise isomorphisms between these different presentations where described in [DF93, FM02].
It has recently emerged that the same trinity of presentations exists in the world of quantum symmetric pairs. A (classical) symmetric pair consists of a semisimple Lie algebra, together with a Lie subalgebra fixed under an involution $\theta$. Their classification is equivalent to the classification of real forms of complex semisimple Lie algebras, and can be formulated combinatorially in terms of Satake diagrams. For example, given an appropriate involution, $(\mathfrak{so}_{n},\mathfrak{sl}_{n})$ is a symmetric pair. However, the situation at the quantum level is more nuanced. The usual quantizations of a Lie algebra and its fixed point subalgebra are typically not compatible. The correct quantization of a symmetric pair was proposed by Letzter [Let99, Let02]. According to her definition, a quantum symmetric pair consists of a quantum group and a coideal subalgebra (rather than a Hopf subalgebra), which specializes to $U(\mathfrak{g}^{\theta})$ in the limit $q\mapsto 1$.
The generalization of quantum symmetric pairs to the Kac–Moody case was carried out by Kolb [Kol14]. In particular, he provided a description of coideal subalgebras by generators and Serre-like relations. The RTT presentation of coideal subalgebras was constructed in special cases by Molev, Ragoucy and Sorba [MRS03], and was preceded by earlier work on twisted Yangians [Ols92]. On the other hand, the Drinfeld-type presentation of affine quantum symmetric pairs is a much more recent development. The first steps in this direction were taken by Baseilhac and Kolb in [BK20], where $q$-root vectors and the relations between them were computed in one of the rank $1$ cases (the $q$-Onsager algebra). Their approach was inspired by Damiani’s work for $\widehat{\mathfrak{sl}}_{2}$ [Dam93], and involved the construction of a new braid group action on the $q$-Onsager algebra, using a computer-based proof. A computer-free proof of the existence of this braid group action was given independently by Terwilliger in [Ter18]. A Drinfeld-type presentation for all simply laced split Satake diagrams was later constructed by Lu and Wang [LW21], using a more general braid group action stemming from a Hall algebra realization of coideal subalgebras [LW23]. A generalization to non-simply laced Satake diagrams was obtained by Zhang [Zha22].
@@ 1.3. Classical motivations
Based on the seminal works of Bao and Wang [BW18a, BW18b], a general philosophy has emerged that most fundamental algebraic, geometric, and categorical constructions from quantum group theory should be generalizable to quantum symmetric pairs. Further advances in this direction were made in [BK19] and [ES18]. Accordingly, Lu and Wang’s discovery of a Drinfeld-type presentation for aKM QSP can be seen as a major step towards generalizing many classical ideas from the theory of quantum affine algebras. Most importantly, these include:
@@
The main goal of this paper is to develop a QSP generalization of i), and provide a foundation for ii). A comprehensive study of $q$-characters for quantum symmetric pairs will be presented in a forthcoming paper with Li [LP24b] (see also [LP24a] and §1.6 below). Some potential implications of our work for iii) are also discussed in §1.7.
@@ 1.4. A shift of perspective
In the foundational works of Letzter [Let99, Let02] and Kolb [Kol14], the coideal structure is of primary importance. However, coideal subalgebras admit a description by generators and relations, and may, accordingly, be regarded as an independent algebraic object (often called an $\imath$quantum group), forgetting the embedding into the quantum group. This point of view is more prevalent in recent literature. It is also rather natural for several reasons. Firstly, $\imath$quantum group constructions are not simple restrictions of the corresponding quantum group constructions. For example, the braid group action from [LWZ23] is not a restriction of Lusztig’s braid group action. Secondly, the proofs are often more technically involved and challenging than their quantum group counterparts, exhibiting many features absent from the latter. Examples include a lack of triangular decompositions and canonical raising and lowering operators, the presence of many rank $1$ cases and parameters, as well as non-obvious integral forms.
@@ 1.5. The main results
What is the relation between the Drinfeld presentation of a quantum affine algebra and Lu and Wang’s Drinfeld-type presentation of a coideal subalgebra?
Perhaps rather surprisingly, we establish that these two presentations are, in fact, mutually compatible in a suitable sense. Before stating our results, let us make a few remarks about the basic assumptions and notation. Let ${\mathfrak{g}}$ be a simple Lie algebra of type $\mathsf{ADE}$. We consider the corresponding untwisted quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$, and its coideal subalgebras $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ of split type. The latter are also often referred to as generalized $q$-Onsager algebras. The coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ depends on two sets of parameters $\mathbf{c}$ and $\mathbf{s}$. The case $\mathbf{s}=0$ is called standard. We use the usual notation for the Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$ (see §8.1 for details). The coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, in its Drinfeld presentation, has two main kinds of generators: $A_{i,k}$ (somewhat analogous to $x_{i,k}^{\pm}$ in $U_{q}\widehat{{\mathfrak{g}}}$ and denoted $B_{i,k}$ in [LW21]) and $\Theta_{i,k}$ (analogous to $\phi_{i,k}^{\pm}$ in $U_{q}\widehat{{\mathfrak{g}}}$). When considering the generators $\Theta_{i,k}$ as a series, we use a slightly different normalization, denoted $\boldsymbol{\grave{\Theta}}_{i}(z)$, than in [LW21]. The reason for it is explained in Remark 4.10.
@@ 1.5.1. Factorization and coproduct theorems
Of particular importance in the Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ is the (almost) commutative subalgebra generated by the elements $\phi_{i,k}^{\pm}$. The spectra of these operators are used to define Drinfeld polynomials and $q$-characters. In a coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, there exists an analogous commutative subalgebra generated by elements $\Theta_{i,k}$. Our key result, which we call the factorization theorem, enables us to express the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ in terms of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$ (see §8.1 for details).
@@
For non-standard parameters $\mathbf{s}\neq 0$, the statement of Theorem 1.1 is slightly more technical, see Theorem 7.5 for details. The restriction to type $\mathsf{A}$ stems from our use of an explicit combinatorial argument related to braid groups (see §1.5.31.5.4 below). Successively more general versions of Theorem 1.1 are proven in Theorems 4.9, 7.5 and Corollary 9.15.
@@
$\Delta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv\boldsymbol{\grave{\Theta}}_{i}(z)\otimes\boldsymbol{\grave{\Theta}}_{i}(z)\quad\mod{(U\otimes U_{+})}[\negthinspace[z]\negthinspace].$
For the proof of Theorem 1.2, see Theorem 7.5 and Corollary 9.15 below. We remark that, since the initial release of this paper, type $\mathsf{BCD}$ versions of the two theorems above have been established in [LP24a].
@@ 1.5.2. Drinfeld rational fractions
Theorem 1.3 (and its close cousins, such as Corollary 5.1 below) is a natural starting point for a $q$-character theory for aKM QSP. For a more precise definition of, as well as a way to compute, Drinfeld rational fractions, see (5.7); for the proof of Theorem 1.3, see Corollary 5.5 below.
@@ 1.5.3. Compatibility of braid group actions
Here $U_{\neq i,+}$ is the subalgebra spanned by elements of positive $Q_{i}$-degree, where $Q_{i}$ is the root sublattice of $Q$ spanned by all positive roots except $\alpha_{i}$. Theorem 1.4 is proven in §9 (see Corollary 9.7). For the details of the notation, see §8.
@@ 1.5.4. A rise in complexity
Our proofs of Theorems 1.11.4 are purely algebraic. However, we would like to point out that their proofs are significantly harder than for some of the corresponding results in the case of quantum affine algebras. For example, the proof of the coproduct formulae for the series $\boldsymbol{\phi}_{i}^{\pm}(z)$ (mod $U_{+}$) relies on relatively straightforward inductions using the Drinfeld presentation. In contrast, the proof for $\boldsymbol{\grave{\Theta}}_{i}(z)$, although also inductive in nature, is much more involved.
The reason for this sharp increase in difficulty can be traced back to the complexity of the Drinfeld presentations themselves. There are two key types of relations which appear both in the original Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ and the Lu–Wang presentation of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. Let us call elements $x_{i,k}^{\pm}$ (resp. $A_{i,k}$) basic, and $h_{i,k},\phi_{i,k}^{\pm}$ (resp. $H_{i,k},\Theta_{i,k}$) diagonal. The first key relation is a commutation relation between basic and diagonal elements, producing further basic elements. In the case of $U_{q}\widehat{{\mathfrak{g}}}$, the RHS of this relation (2.7) involves only one term, while in the relation (2.17) for $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ two terms are present. The second key relation is a commutation relation between basic elements, producing diagonal elements. Again, this relation (2.9) is much easier for $U_{q}\widehat{{\mathfrak{g}}}$, involving only a simple commutator on the LHS and typically one term on the RHS. In contrast, the corresponding relation (2.18) for $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ involves two $q$-commutators on the LHS and typically two terms on the RHS.
@@ 1.6. Application to $q$-characters
The results of this paper are sufficient to formulate a notion of $q$-characters for quantum symmetric pairs and show that such characters are compatible, in an appropriate sense, with the usual $q$-characters for quantum affine algebras. A detailed exposition, for all the classical types, can be found in [LP24a]. Let us summarize here the main ideas.
@@
Therefore, we propose to define boundary $q$-characters directly via the Lu–Wang presentation. More precisely, we consider the generalized eigenspace decomposition of a finite dimensional representation with respect to the action of the operators $\Theta_{i,m}$, and let boundary $q$-characters encode the multiplicities of such eigenspaces. This is equivalent to taking the trace of a certain operator in the completion ${\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, yielding a map $\chi^{\imath}_{q}\colon\operatorname{Rep}\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})\to{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$. Applying Theorems (1.1)–(1.3), one can show that $\chi^{\imath}_{q}$ is compatible with the usual $q$-character map ([LP24a, Corollary 9.6]). More precisely, the diagram

### v6

@@ Abstract.
We formulate a precise connection between the new Drinfeld presentation of a quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$ and the new Drinfeld presentation of affine coideal subalgebras of split type recently discovered by Lu and Wang. In particular, we establish a “factorization formula”, expressing the commuting “Drinfeld–Cartan”-type operators $\Theta_{i,k}$ in the coideal subalgebra in terms of the corresponding Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$. We study the spectra of these operators on finite dimensional representations, and describe them in terms of rational functions with an extra symmetry. These results can be seen as the starting point of a $q$-character theory for affine Kac–Moody quantum symmetric pairs. Additionally, we prove a compatibility result linking Lusztig’s and the Lu-Wang-Zhang braid group actions.
@@ 1.1. A trinity of presentations
It is well known that quantum affine algebras $U_{q}\widehat{{\mathfrak{g}}}$ admit three distinct presentations: the Kac–Moody (also: Drinfeld–Jimbo, Serre) presentation [Jim85, Dri87b], the new Drinfeld (or loop) presentation [Dri87a, Bec94, Dam12, Dam15] and the FRT (also: RTT) presentation [RSTS90]. The first of these can be seen as a quantization of the usual presentation of an affine Kac–Moody Lie algebra, while the second quantizes its construction as a central extension of a loop Lie algebra. Precise isomorphisms between these different presentations were described in [DF93, FM02].
It has recently emerged that the same trinity of presentations exists in the world of quantum symmetric pairs. A (classical) symmetric pair $({\mathfrak{g}},{\mathfrak{g}}^{\theta})$ consists of a semisimple Lie algebra ${\mathfrak{g}}$, together with a Lie subalgebra fixed under an involution $\theta$. We note that ${\mathfrak{g}}^{\theta}$ is reductive. The classification of symmetric pairs is equivalent to the classification of real forms of complex semisimple Lie algebras, and can be formulated combinatorially in terms of Satake diagrams. For example, given an appropriate involution, $(\mathfrak{sl}_{n},\mathfrak{so}_{n})$ is a symmetric pair. However, the situation at the quantum level is more nuanced. The Drinfeld–Jimbo quantizations of a Lie algebra and its fixed point subalgebra are typically not compatible. The correct quantization of a symmetric pair was proposed by Letzter [Let99, Let02]. According to her definition, a quantum symmetric pair consists of a quantum group and a coideal subalgebra (rather than a Hopf subalgebra), which specializes to $U(\mathfrak{g}^{\theta})$ in the limit $q\mapsto 1$.
The generalization of quantum symmetric pairs to the Kac–Moody case was carried out by Kolb [Kol14]. In particular, he provided a description of coideal subalgebras by generators and Serre-like relations. The RTT presentation of coideal subalgebras was constructed in special cases by Molev, Ragoucy and Sorba [MRS03], and was preceded by earlier work on twisted Yangians [Ols92]. On the other hand, the Drinfeld-type presentation of affine quantum symmetric pairs is a much more recent development. The first steps in this direction were taken by Baseilhac and Kolb in [BK20], where $q$-root vectors and the relations between them were computed in one of the rank $1$ cases (the $q$-Onsager algebra). Their approach was inspired by Damiani’s work for $\widehat{\mathfrak{sl}}_{2}$ [Dam93], and involved the construction of a new braid group action on the $q$-Onsager algebra, using a computer-based proof. A computer-free proof of the existence of this braid group action was given independently by Terwilliger in [Ter18]. A Drinfeld-type presentation for all simply laced split Satake diagrams was later constructed by Lu and Wang [LW21], using a more general braid group action stemming from a Hall algebra realization of coideal subalgebras [LW23]. A generalization to non simply-laced split Satake diagrams was obtained by Zhang [Zha22].
@@ 1.3. Classical motivations
In the seminal works of Bao and Wang [BW18a, BW18b], a general philosophy was proposed that most fundamental algebraic, geometric, and categorical constructions from quantum group theory should be generalizable to quantum symmetric pairs. Around the same time, advances in this direction were also made in the influential papers [BK19] and [ES18]. Accordingly, Lu and Wang’s discovery of a Drinfeld-type presentation for aKM QSP can be seen as a major step towards generalizing many classical ideas from the theory of quantum affine algebras. Most importantly, these include:
@@
The main goal of this paper is to develop a QSP generalization of i), and provide a foundation for ii). A comprehensive study of $q$-characters for quantum symmetric pairs will be presented in a forthcoming paper with Li [LP25b] (see also [LP25a] and §1.6 below). Some potential implications of our work for iii) are also discussed in §1.7.
@@ 1.4. A shift of perspective
In the foundational works of Letzter [Let99, Let02] and Kolb [Kol14], the coideal structure is of primary importance. However, QSP coideal subalgebras admit a description by generators and relations, and may, accordingly, be regarded as an independent algebraic object (often called an $\imath$quantum group), forgetting the embedding into the quantum group. This point of view is more prevalent in recent literature. It is also rather natural for several reasons. Firstly, $\imath$quantum group constructions are not simple restrictions of the corresponding quantum group constructions. For example, the braid group action from [LWZ23] is not a restriction of Lusztig’s braid group action. Secondly, the proofs are often more technically involved and challenging than their quantum group counterparts, exhibiting many features absent from the latter. Examples include a lack of triangular decompositions and canonical raising and lowering operators, the presence of many rank $1$ cases and parameters, as well as non-obvious integral forms.
@@ 1.5. The main results
What is the relation between the Drinfeld presentation of a quantum affine algebra and Lu and Wang’s Drinfeld-type presentation of a QSP coideal subalgebra?
Remarkably, we establish that these two presentations are, in fact, mutually compatible in a suitable sense. Before stating our results, let us make a few remarks about the basic assumptions and notations. Let ${\mathfrak{g}}$ be a simple Lie algebra of type $\mathsf{ADE}$. We consider the corresponding untwisted quantum affine algebra $U_{q}\widehat{{\mathfrak{g}}}$, and its coideal subalgebras $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ of split type. The latter are also often referred to as generalized $q$-Onsager algebras. The coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ depends on two sets of parameters $\mathbf{c}$ and $\mathbf{s}$ [^1] In this introduction, in order to simplify notation, we will suppress the dependence on the parameters. In other words, $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ refers here to a QSP coideal subalgebra with chosen $\mathbf{c}$ and $\mathbf{s}$. Later, in §2 and §8, $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ will refer to the universal version of this algebra, where the parameters $\mathbf{c}$ are replaced by central elements. The coideal subalgebras with fixed parameters will then be denoted as $\mathcal{O}_{q}^{\mathbf{c},\mathbf{s}}(\widehat{{\mathfrak{g}}})$.. The case $\mathbf{s}=0$ is called standard. We use the usual notation for the Drinfeld generators of $U_{q}\widehat{{\mathfrak{g}}}$, with “raising operators” $x_{i,k}^{+}$, “lowering operators” $x_{i,k}^{-}$ and “Cartan operators” $h_{i,k},\phi^{\pm}_{i,k}$ (see §8.1 for details). The coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, in its Drinfeld presentation, has two main kinds of generators: $A_{i,k}$ (somewhat analogous to $x_{i,k}^{\pm}$ in $U_{q}\widehat{{\mathfrak{g}}}$ and denoted $B_{i,k}$ in [LW21]) and $\Theta_{i,k}$ (analogous to $\phi_{i,k}^{\pm}$ in $U_{q}\widehat{{\mathfrak{g}}}$). When considering the generators $\Theta_{i,k}$ as a series, we use a slightly different normalization, denoted $\boldsymbol{\grave{\Theta}}_{i}(z)$, than in [LW21]. The reason for it is explained in Remark 4.10.
@@ 1.5.1. Factorization and coproduct theorems
Of particular importance in the Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ is the (almost) commutative subalgebra generated by the elements $\phi_{i,k}^{\pm}$. The spectra of these operators are used to define Drinfeld polynomials and $q$-characters. In a coideal subalgebra $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$, there exists an analogous commutative subalgebra generated by elements $\Theta_{i,k}$. Our key result, which we call the factorization theorem, enables us to express the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ in terms of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$, modulo the “Drinfeld positive half” of $U_{q}\widehat{{\mathfrak{g}}}$, i.e., the subalgebra spanned by elements of positive degree in the Drinfeld gradation, where $\operatorname{deg}^{\mathrm{Dr}}x^{\pm}_{i,k}=\pm\alpha_{i},\ \operatorname{deg}^{\mathrm{Dr}}h_{i,k}=0$ (see §8.1 for details).
@@
Here $C$ is a scalar depending on the choice of parameters $\mathbf{c}$. For non-standard parameters $\mathbf{s}\neq 0$, the statement of Theorem 1.1 is slightly more technical, see Theorem 7.5 for details. The restriction to type $\mathsf{A}$ stems from our use of an explicit combinatorial argument related to braid groups (see §1.5.31.5.4 below). Successively more general versions of Theorem 1.1 are proven in Theorem 4.9 (rank one, standard parameters), Theorem 7.5 (rank one, nonstandard parameters) and Corollary 9.16 (arbitrary rank).
@@
$\Delta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv\boldsymbol{\grave{\Theta}}_{i}(z)\otimes\boldsymbol{\grave{\Theta}}_{i}(z)\quad\mod{(\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})\otimes U_{+})}[\negthinspace[z]\negthinspace].$
For the proof of Theorem 1.2, also see Theorem 7.5 and Corollary 9.16 below. We remark that, since the initial release of this paper, type $\mathsf{BCD}$ versions of the two theorems above have been established in [LP25a].
@@ 1.5.2. Drinfeld rational fractions
Theorem 1.3 (and its close cousins, such as Corollary 5.1 below) is a natural starting point for a $q$-character theory for aKM QSP. For a more precise definition of, as well as a way to compute, Drinfeld rational fractions, see (5.7); for the proof of Theorem 1.3, see Theorem 5.5 below.
@@ 1.5.3. Compatibility of braid group actions
Here $U_{\neq i,+}$ is the subalgebra spanned by elements of positive $Q_{i}$-degree, where $Q_{i}$ is the root sublattice of $Q$ spanned by all positive roots except $\alpha_{i}$. Theorem 1.4 is proven in §9 (see Corollary 9.8). For the details of the notation, see §8. The generalization of Theorem 1.4 to $\mathsf{BCD}$ types can be found in [LP25a].
@@ 1.5.4. A rise in complexity
Our proofs of Theorems 1.11.4 are purely algebraic. However, they are significantly harder than the proofs of the corresponding results for quantum affine algebras. The reason for this sharp increase in difficulty can be traced back to the complexity of the Drinfeld presentations themselves. There are two key types of relations which appear both in the original Drinfeld presentation of $U_{q}\widehat{{\mathfrak{g}}}$ and the Lu–Wang presentation of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$. Let us call elements $x_{i,k}^{\pm}$ (resp. $A_{i,k}$) basic, and $h_{i,k},\phi_{i,k}^{\pm}$ (resp. $H_{i,k},\Theta_{i,k}$) diagonal. The first key relation is a commutation relation between basic and diagonal elements, producing further basic elements. In the case of $U_{q}\widehat{{\mathfrak{g}}}$, the RHS of this relation (2.7) involves only one term, while in the relation (2.18) for $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ two terms are present. The second key relation is a commutation relation between basic elements, producing diagonal elements. Again, this relation (2.9) is much easier for $U_{q}\widehat{{\mathfrak{g}}}$, involving only a simple commutator on the LHS and typically one term on the RHS. In contrast, the corresponding relation (2.19) for $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ involves two $q$-commutators on the LHS and typically two terms on the RHS.
@@ 1.6. Application to $q$-characters
The results of this paper are sufficient to formulate a notion of $q$-characters for quantum symmetric pairs and show that such characters are compatible, in an appropriate sense, with the usual $q$-characters for quantum affine algebras. A detailed exposition, for all the classical types, can be found in [LP25a]. Let us summarize here the main ideas.
@@
Therefore, we propose to define boundary $q$-characters directly via the Lu–Wang presentation. More precisely, we consider the generalized eigenspace decomposition of a finite dimensional representation with respect to the action of the operators $\Theta_{i,m}$, and let boundary $q$-characters encode the multiplicities of such eigenspaces. This is equivalent to taking the trace of a certain operator in the completion ${\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, yielding a map $\chi^{\imath}_{q}\colon\operatorname{Rep}\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})\to{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$. Applying Theorems (1.1)–(1.3), one can show that $\chi^{\imath}_{q}$ is compatible with the usual $q$-character map ([LP25a, Corollary 9.3]). More precisely, the diagram